# Federated Unlearning: A Survey on Methods, Design Guidelines, and Evaluation Metrics

### v1

@@ V-A Client Unlearning
Most of the methods proposed for client unlearning can be seen as fast retraining strategies, which design specific ways to retain as much knowledge as possible from the ongoing training while selectively overwriting the influence of forgetting data.
@@
Gradient Modification. In [49], Halimi et al. introduce a method to perform unlearning directly on the target client. The method employs projected gradient ascent to maximize the loss from a reference model, defined as the average of the models of the remaining clients. A $\ell_{2}$-norm sphere with $\delta$ radius around the reference model that constrains the unbounded loss. To improve the performance of the unlearned model, the remaining clients run additional rounds of FL, achieving good performance in a few rounds. Similarly, Li et al. [61] propose a subspace-based FU method (SFU) that leverages gradient ascent, constrained by the orthogonal space of input gradient spaces formed by other clients, to eliminate the target client’s contribution. SFU does not require storing intermediate updates on the clients or the server. The process includes three steps: (1) each client, except the target one, creates a representation matrix using a part of its local data and sends it to the server. A representation matrix for each layer in a neural network is formed by combining outputs from the forward pass of random samples through the network, with each column representing the output for a different sample. The privacy of the representation matrix is protected through a DP algorithm. Random factors are added to each client’s layer representation to prevent the leakage of individual client data information. This addition does not affect the subspace search process due to the orthogonal nature of the matrix. (2) The target client executes some round of gradient ascent and sends the updated gradient to the server. (3) The server then performs Singular Value Decomposition (SVD) [78] on the set of representation matrices to get the set of input gradient subspace. It gets the orthogonal projection of the target gradient for this subspace and removes the contribution of the target client by updating the global model.
@@
Other Approaches. RevFRF [67] is a framework designed for federated Random Forests (RFs) that supports secure participant revocation. The authors delve into the revocation process from two distinctive perspectives: (1) ensuring the removal of targeted data from the RF, and (2) preempting the potential collusion of removed clients with the server to illicitly continue utilizing the outdated RF. To achieve the first level of revocation, the RevFRF traverses the random decision trees (RDTs) within the trained RF and prunes all splits contributed by the participant seeking to be forgotten. Subsequently, the RF undergoes a reconstruction, wherein only the remaining RDTs are considered. Due to their isolation from each other, the rebuilding of one RDT does not affect the others. It is crucial to note, however, that this approach is not easily extendable to other ML models, and the worst-case scenario coincides with the reconstruction of the model from scratch. The second level of revocation involves executing additional computations to refresh the revoked splits with random values, rendering them inaccessible to any potentially colluding server.
@@ V-C Sample Unlearning
Dhasade et al. [69] introduce QuickDrop, an FU method that leverages Dataset Distillation (DD) to perform unlearning and recovery. DD is a technique to condense a large training dataset into a compact and smaller synthetic dataset [82]. However, the production of good-quality synthetic data requires many local update steps. The clients reuse the gradient updates computed during FL training to reduce the computational overhead. When an unlearning request comes, each client performs Stochastic Gradient Ascent (SGA) on their local distilled dataset. While, in the recovery phase, the network then executes recovery rounds, during which they also use the distilled data that are not extracted from the samples to forget. It is worth noting that the use of distilled data slightly decreases the performance. To avoid decreasing performance, in the recovery phase, the authors include a few original samples in the distilled datasets that nullify potential performance drops.
@@
Other Approaches. FL can be also employed with alternative data representation such as a Knowledge Graph (KG), a structured knowledge base of real-world entities and their relations in terms of triplets. KGs have also been adopted to advance representation learning techniques, i.e., KG embeddings combine entities and their relations into a unified semantic space [84]. FL can be integrated with KG embeddings, giving birth to a novel paradigm namely Federated KG embedding learning [85]. In this direction, Zhu et al. [75] propose FedLU, an FL framework for heterogeneous KG embedding. FedLU uses an unlearning method to erase specific knowledge from local embeddings. On the one hand, local knowledge from clients is erased through an interference phase based on hard and soft confusions, which reflect the distance between scores of the triplet in the forgetting set and its negative samples, i.e., the set of triplets whose intersection with the local KG is empty. On the other hand, to limit the decrease in model performance caused by the interference is recovered via knowledge distillation providing the local client KG without the forgetting set.
@@ VI Distilled Lessons Learned
    Motivation: FU schemes are proposed as valuable mechanisms to preserve privacy and/or enhance security. From the privacy perspective, FU can guarantee the right to be forgotten. On the security side, it is a technique for mitigating the impact of malicious clients by eliminating their contributions, i.e., poisoned data and backdoors.
@@
    Empirical Evidences: The contribution of a specific client $u$ has a tangible influence on the global model, which outputs significantly more accurate predictions when presented with client $u$’s samples if client $u$ has been included in the pool of clients. Also, the participation of client $u$ translates to evident spikes of enhanced accuracy (and reduced loss) on client $u$’s train data for that round. Natural forgetting, i.e. just excluding client $u$ from the federation in the hope of catastrophic forgetting to rapidly remove its contribution, is not a viable option. Indeed, empirical evidences show that the impact of a client participation fades away very slowly when data is homogeneous, and remains concretely noticeable for a large number of subsequent rounds. It is worth noting that in the particular case of a large-scale federation with low participation rate of clients and non-IID local datasets, natural forgetting can help the unlearning process.
@@
    Requirements and Metrics: As we have detailed from the requirements in Sec. IV-A, an unlearning algorithm should be more efficient than a retraining alternative, and the (fine-tuned) unlearned global model should achieve performance of generalization comparable to the original model. At the same time, with effective unlearning, the unlearned global model should not respond suspiciously confidently when presented with forgetting data. Then, how to verify those requirements?
@@
    Methods: The majority of the presented solutions center around the complete removal of a specific client’s contributions, i.e., client unlearning. However, the algorithms designed for a particular FU objective may have broader applicability. Specifically, client unlearning coincides with sample unlearning when the samples marked for forgetting are exclusively provided by a singular client and encompass all of their data. Furthermore, the concept of client unlearning can be viewed as a specialized instance of class unlearning, applicable when only a single client contributes to the category targeted for erasure. Similarly, these considerations can be extended to sample unlearning, wherein the subset marked for forgetting corresponds to a class to be unlearned.
@@ VII-A Standardization
However, the lack of any standardized datasets and metrics in MU and consequently in FU demonstrate that a significant amount of work needs to be done in this direction. While some efforts have been made in the context of FL (e.g. LEAF [86], TensorFlow Federated datasets [87] and Flower datasets [88, 89]), extending such standardization to MU and FU is an area that requires substantial attention and development to foster more reliable and comparable research outcomes.
@@ VII-C Multiple Unlearning Request
Although in real-world scenarios clients may continually request multiple, sequential unlearning requests, almost all the existing FU methods do not take this condition into account [69].
@@ VII-D Feature Unlearning
Despite the potential of feature unlearning is still an open challenge, as supported by the very limited body of research. Deng et al. [91] contribute to addressing this gap by presenting a method based on logistic regression for vertical FU. In the context of VFL, participants collaboratively share a common set of data samples while maintaining distinct sets of features [92]. The proposed unlearning algorithm subtracts the updates from the target client, contributing to the refinement of the global model. To enhance both efficiency and accuracy, constraints are imposed on intermediate updates, thereby minimizing the sum of intermediate parameters. This optimization strategy results in a reduced impact on the sum of intermediate parameters for the remaining clients when the target client’s contributions are removed. To generate the unlearned model, the server must retain the intermediate parameters submitted by each client during the overall training. When a target client requests unlearning of their data, the server retrieves the intermediate updates from that client in the last round, negates these updates, and distributes them to the remaining clients. Subsequently, each client utilizes these negated updates to adjust their local model parameters
@@ VIII Conclusions
This paper presents a comprehensive survey on FU, with a particular focus on technical design/implementation guidelines and practical usage scenarios. Firstly, we introduce the background concepts needed to understand the FU research area, by including the formalization of FU objectives and the clarifications of the key distinctions from traditional MU. To provide concrete insights into FU and its implications, we conduct a series of experiments by using various datasets. Subsequently, we offer valuable guidelines for designing and implementing FU algorithms, encompassing an in-depth analysis of the most employed FU evaluation metrics. Moreover, the paper offers a technically detailed review of the existing literature, by categorizing the proposed solutions according to a novel taxonomy based on FU objectives and metrics employed. Finally, we discuss open challenges and identify the most promising directions for future research work in the field.
Appendix A Experimental Setup
Figure 6: The label distribution across clients, i.e. the per-label number of local examples in each client’s dataset.
@@ Appendix A Experimental Setup
This section describes the experimental setting for the results reported in Table III, Fig. 4, Fig. 7, Fig. 5, and Fig. 8. We implemented the simulations with Flower [88] and TensorFlow. Our code is publicly available at www.github.com/alessiomora/unlearning_fl.
A-A Model Architecture and Hyperparameters
@@
We used a visual transformer, i.e. MiT-B0 [33], with approximately 3.6M parameters, initialized from a pre-trained model checkpoint trained on ImageNet-1k (69.27% accuracy on test data). We adapted the one-layer classification head to the specific task, initializing such a layer from scratch. We employed the AdamW optimizer with a client learning rate of 3e-4, with a round-wise exponential decay of 0.998, 5 local epochs, batch size of 32, and weight decay regularization of 1e-3.
A-B Datasets
@@
We performed the experiments on three datasets: CIFAR-100 [100], Caltech-2011 (birds) [101], and FGVC-Aircraft (aircraft) [102]. We provided two settings for the data partitioning among clients of each dataset: identically and independently distributed (IID) data, i.e. each client holds approximately the same number of per-class examples, and non-IID simulated via distribution-based label skew following the method from [37]. For the latter setting, we tuned the label skew using a concentration parameter of $\alpha=0.1$ to rule the Dirichlet distribution, with Fig. 6 showing the label distribution across clients. The same data example is not repeated among multiple clients.
@@
CIFAR-100. CIFAR-100 consists of 60,000 examples of 32x32 color images – 50,000 for training and 10,000 for testing – belonging to 100 classes. To match the transformer models’ input size, we resized the images to a resolution of 224x224 pixels; we also preprocessed the training images with random crop and horizontal flip layers. For the experiments reported in Fig. 4, in Fig. 7 and in Table III, we partitioned the training set to simulate 100 clients in the federation; we set 100 as the number of clients so that, in the IID setting, each client can have at least five per-class (unique) examples. At each round, 10 clients out of 100 were randomly selected to participate. For the experiments on natural forgetting, reported in Fig. 5 and 8, we partitioned the training set to simulate 500 clients in the federation, with 20 clients selected per round.
@@
Caltech-2011. Caltech 2011 (birds) consists of 11,788 examples of color images – 5,994 for training and 5,794 for testing – belonging to 200 classes. To match the input size of the transformer models we resized the images to a resolution of 224x224 pixels; we also preprocessed the training images with random crop and horizontal flip layers, similarly to the work in [103]. We partitioned the training set to simulate 29 clients in the federation; we set 29 as the number of clients so that, in the IID setting, each client can have at least one per-class (unique) example. At each round, 5 clients out of 29 were randomly selected to participate.
@@
FGVC-Aircraft. The FGVC-Aircraft dataset contains 10,000 images of aircraft – 6,667 for training and validation and 3,333 for testing. The aircraft labels are organized in a four-level hierarchy, i.e., model, variant, family, and manufacturer (from finer to coarser). We consider a classification at the variant level (e.g. Boeing 737-700). A variant collapses all the models that are visually indistinguishable into one class. The dataset contains 100 images for each of the 100 different aircraft model variants. We removed the copyright banner from the images by cutting off the bottom 20 pixels in height. To match the input size of the transformer models we resized the images to a resolution of 224x224 pixels; we also preprocessed the training images with random crop and horizontal flip layers, similarly to the work in [103]. We partitioned the training set to simulate 65 clients in the federation; we set 65 as the number of clients so that, in the IID setting, each client can have at least one per-class (unique) example. At each round, 7 clients out of 65 were randomly selected to participate.
Table VI summarizes the dataset-specific configurations.
A-C Metrics
We tracked three main metrics: accuracy, loss, and KL divergence. Accuracy and loss are measured either on test data or on client $u$’s train data. The third metric calculates the KL divergence between the output distribution of the model on client $u$’s data and a uniformly distributed output probability. This metric should work as a proxy for the meaningfulness of model predictions (the greater the divergence the more meaningful the predictions).
Figure 7: The global model’s loss on test data (solid line) versus the global model’s loss on client $u$’s train data (dashed line) across many FL rounds. Grey bars indicate that client $u$ has participated in that round. The upper row of charts considers IID data distribution among clients. The bottom row of charts considers non-IID data distribution among clients, simulated via distribution-based labels skew (with concentration parameter $\alpha$=0.1).
Figure 8: KL divergence, loss and accuracy tracked for 100 rounds. 500 clients (4% participation rate). For the red line, client $u$ participates only during the first 50 rounds. For the gray line, client $u$ never participates.

### v3

@@ V-A Client Unlearning
Most methods proposed for client unlearning can be seen as fast retraining strategies, which design specific ways to retain as much knowledge as possible from the ongoing training while selectively overwriting the influence of forgetting data.
@@
Gradient Modification. In [50], Halimi et al. introduce a method to perform unlearning directly on the target client. The method employs projected gradient ascent to maximize the loss from a reference model, defined as the average of the models of the remaining clients. A $\ell_{2}$-norm sphere with $\delta$ radius around the reference model is used to constrain the unbounded loss. To improve the performance of the unlearned model, the remaining clients run additional rounds of FL, achieving good performance in a few rounds. Similarly, Li et al. [62] propose a subspace-based FU method (SFU) that leverages gradient ascent, constrained by the orthogonal space of input gradient spaces formed by other clients, to eliminate the target client’s contribution. SFU does not require storing intermediate updates on the clients or the server. The process includes three steps: (1) each client, except the target one, creates a representation matrix using a part of its local data and sends it to the server. A representation matrix for each layer in a neural network is formed by combining outputs from the forward pass of random samples through the network, with each column representing the output for a different sample. The privacy of the representation matrix is protected through a DP algorithm. Random factors are added to each client’s layer representation to prevent the leakage of individual client data information. This addition does not affect the subspace search process due to the orthogonal nature of the matrix. (2) The target client executes some round of gradient ascent and sends the updated gradient to the server. (3) The server then performs Singular Value Decomposition (SVD) [80] on the set of representation matrices to get the set of input gradient subspace. It gets the orthogonal projection of the target gradient for this subspace and removes the contribution of the target client by updating the global model.
@@
Other Approaches. Tao et al. [70] propose a framework for exact FU that leverages the concept of total variation (TV) stability, which measures the discrepancy between two distributions. A learning algorithm is $\rho$-TV-stable if the model produced using the entire dataset and the model produced using a subset of it differ by at most $\rho$. The authors introduce FATS, a novel algorithm for fast retraining based on FedAvg and proved to be $\rho$-TV-stable.
RevFRF [68] is a framework designed for federated Random Forests (RFs) that supports secure participant revocation. The authors delve into the revocation process from two distinctive perspectives: (1) ensuring the removal of targeted data from the RF, and (2) preempting the potential collusion of removed clients with the server to illicitly continue utilizing the outdated RF. To achieve the first level of revocation, the RevFRF traverses the Random Decision Trees (RDTs) within the trained RF and prunes all splits contributed by the participant seeking to be forgotten. Subsequently, the RF undergoes a reconstruction, wherein only the remaining RDTs are considered. Due to their isolation from each other, the rebuilding of one RDT does not affect the others. It is crucial to note, however, that this approach is not easily extendable to other ML models, and the worst-case scenario coincides with the reconstruction of the model from scratch. The second level of revocation involves executing additional computations to refresh the revoked splits with random values, rendering them inaccessible to any potentially colluding server.
@@ V-C Sample Unlearning
Dhasade et al. [71] introduce QuickDrop, an FU method that leverages Dataset Distillation (DD) to perform unlearning and recovery. DD condenses a large training dataset into a compact and smaller synthetic dataset [84]. However, generating good-quality synthetic data requires many local update steps. The clients reuse the gradient updates computed during FL training to reduce the computational overhead. When an unlearning request comes, each client performs Stochastic Gradient Ascent (SGA) on their local distilled dataset. While in the recovery phase, the network then executes recovery rounds, during which they also use the distilled data that are not extracted from the samples to forget. It is worth noting that the use of distilled data slightly decreases the performance. To avoid decreasing performance, in the recovery phase, the authors include a few original samples in the distilled datasets that nullify potential performance drops.
@@
Other Approaches. The FATS algorithm, presented in [70], can be also applied for sample unlearning by starting the retraining from the iteration preceding the inclusion of that specific sample.
FL can be also employed with alternative data representation such as a Knowledge Graph (KG), a structured knowledge base of real-world entities and their relations in terms of triplets. KGs have also been adopted to advance representation learning techniques, i.e., KG embeddings combine entities and their relations into a unified semantic space [86]. FL can be integrated with KG embeddings, giving birth to a novel paradigm namely Federated KG embedding learning [87]. In this direction, Zhu et al. [77] propose FedLU, an FL framework for heterogeneous KG embedding. FedLU uses an unlearning method to erase specific knowledge from local embeddings. On the one hand, local knowledge from clients is erased through an interference phase based on hard and soft confusions, which reflect the distance between scores of the triplet in the forgetting set and its negative samples, i.e., the set of triplets whose intersection with the local KG is empty. On the other hand, to limit the decrease in model performance caused by the interference is recovered via knowledge distillation providing the local client KG without the forgetting set.
@@ VI Distilled Lessons Learned
    Motivation. FU schemes are proposed as valuable mechanisms to preserve privacy and/or enhance security. From the privacy perspective, FU can guarantee the right to be forgotten. On the security side, it is a technique for mitigating the impact of malicious clients by eliminating their contributions, i.e., poisoned data and backdoors.
@@
    Empirical Evidences. The contribution of a specific client $u$ has a tangible influence on the global model, which outputs significantly more accurate predictions when presented with client $u$’s samples if client $u$ has been included in the pool of clients. Also, the participation of client $u$ translates to evident spikes of enhanced accuracy (and reduced loss) on client $u$’s train data for that round. Natural forgetting, i.e., just excluding client $u$ from the federation in the hope of catastrophic forgetting to rapidly remove its contribution, is not a viable option. Indeed, empirical evidences show that the impact of client participation fades away very slowly when data is homogeneous, and remains concretely noticeable for a large number of subsequent rounds. It is worth noting that in the particular case of a large-scale federation with a low participation rate of clients and non-IID local datasets, natural forgetting can help the unlearning process.
@@
    Requirements and Metrics. As we have detailed from the requirements in Sec. IV-A, an unlearning algorithm should be more efficient than a retraining alternative, and the (fine-tuned) unlearned global model should achieve performance of generalization comparable to the original model. At the same time, with effective unlearning, the unlearned global model should not respond suspiciously confidently when presented with forgetting data. Then, how to verify those requirements?
@@
    Methods. The majority of the presented solutions center around the complete removal of a specific client’s contributions, i.e., client unlearning. However, the algorithms designed for a particular FU objective may have broader applicability. Specifically, client unlearning coincides with sample unlearning when the samples marked for forgetting are exclusively provided by a singular client and encompass all of their data. Furthermore, the concept of client unlearning can be viewed as a specialized instance of class unlearning, applicable when only a single client contributes to the category targeted for erasure. Similarly, these considerations can be extended to sample unlearning, wherein the subset marked for forgetting corresponds to a class to be unlearned.
@@ VII-A Standardization
However, the lack of any standardized datasets and metrics in MU and consequently in FU demonstrate that a significant amount of work needs to be done in this direction. While some efforts have been made in the context of FL (e.g., LEAF [88], TensorFlow Federated datasets [89] and Flower datasets [90, 91]), extending such standardization to MU and FU is an area that requires substantial attention and development to foster more reliable and comparable research outcomes.
@@ VII-C Multiple Unlearning Request
Although in real-world scenarios clients may continually request multiple, sequential unlearning requests, almost all the existing FU methods overlook this condition [71].
@@ VII-D Feature Unlearning
Despite the potential of feature unlearning is still an open challenge, as supported by the very limited body of research. Deng et al. [93] contribute to addressing this gap by presenting a method based on logistic regression for vertical FU. In the context of VFL, participants collaboratively share a common set of data samples while maintaining distinct sets of features [94]. The proposed unlearning algorithm subtracts the updates from the target client, contributing to the refinement of the global model. To enhance both efficiency and accuracy, constraints are imposed on intermediate updates, thereby minimizing the sum of intermediate parameters. This optimization strategy results in a reduced impact on the sum of intermediate parameters for the remaining clients when the target client’s contributions are removed. To generate the unlearned model, the server must retain the intermediate parameters submitted by each client during the overall training. When a target client requests unlearning of their data, the server retrieves the intermediate updates from that client in the last round, negates these updates, and distributes them to the remaining clients. Subsequently, each client utilizes these negated updates to adjust their local model parameters.
@@ VIII Conclusions
This paper presents a comprehensive survey on FU, with a particular focus on technical design/implementation guidelines and practical usage scenarios. Firstly, we introduce the background concepts needed to understand the FU research area, by including the formalization of FU objectives and the clarifications of the key distinctions from traditional MU. To provide concrete insights into FU and its implications, we conduct a series of experiments by using various datasets. Subsequently, we offer valuable guidelines for designing and implementing FU algorithms, encompassing an in-depth analysis of the most employed FU evaluation metrics. Moreover, the paper offers a technically detailed review of the existing literature, by categorizing the proposed solutions according to a novel taxonomy based on the target FU objectives and the employed metrics. Finally, we contribute to the technical discussion in the field by identifying the most relevant open challenges and the most promising directions for future research work.
Appendix - Experimental Setup
Figure 6: KL divergence and accuracy for 100 rounds. 500 clients (4% participation rate). Red line: client $u$ participates only during the first 50 rounds. Gray line: client $u$ never participates.
@@
This section describes the experimental setting for the results reported in Table III, Figure 4, Figure 5, and Figure 6. We implemented the simulations with Flower [90] and TensorFlow. Our code is publicly available at www.github.com/alessiomora/unlearning_fl.
VIII-A Model Architecture and Hyperparameters
@@
We used a visual transformer, i.e., MiT-B0 [35], with approximately 3.6M parameters, initialized from a pre-trained model checkpoint trained on ImageNet-1k (69.27% accuracy on test data). We adapted the one-layer classification head to this task, initializing such a layer from scratch. We employed the AdamW optimizer with a client learning rate of 3e-4, with a round-wise exponential decay of 0.998, 5 local epochs, batch size of 32, and weight decay regularization of 1e-3.
VIII-B Datasets
@@
We performed experiments on three datasets: CIFAR-100 [102], Caltech-2011 (birds) [103], and FGVC-Aircraft (aircraft) [104]. We provided two settings for the data partitioning among clients: identically and independently distributed (IID) data, i.e., each client holds approximately the same number of per-class examples, and non-IID simulated via distribution-based label skew following the method from [39]. For the latter setting, we tuned the label skew using a concentration parameter of $\alpha=0.1$ to rule the Dirichlet distribution.The same data example is not repeated among multiple clients.
@@
CIFAR-100. CIFAR-100 consists of 60,000 examples of 32x32 color images – 50,000 for training and 10,000 for testing – belonging to 100 classes. To match the transformer models’ input size, we resized the images to a resolution of 224x224 pixels; we also preprocessed the training images with random crop and horizontal flip layers. For the experiments reported in Figure 4 and Table III, we partitioned the training set to simulate 100 clients in the federation; we set 100 as the number of clients so that, in the IID setting, each client can have at least five per-class (unique) examples. At each round, 10 clients out of 100 were randomly selected to participate. For the experiments on natural forgetting, reported in Figure 5 and 6, we partitioned the training set to simulate 500 clients in the federation, with 20 clients selected per round.
@@
Caltech-2011. Caltech 2011 (birds) consists of 11,788 examples of color images – 5,994 for training and 5,794 for testing – belonging to 200 classes. To match the input size of the transformer models we resized the images to a resolution of 224x224 pixels; we also preprocessed the training images with random crop and horizontal flip layers, similarly to the work in [105]. We partitioned the training set to simulate 29 clients; we set 29 as the number of clients so that, in the IID setting, each client can have at least one per-class (unique) example. At each round, 5 clients out of 29 were selected.
@@
FGVC-Aircraft. The FGVC-Aircraft dataset contains 10,000 images of aircraft – 6,667 for training and validation and 3,333 for testing. The aircraft labels are organized in a four-level hierarchy. We consider a classification at the variant level (e.g., Boeing 737-700). A variant collapses all the models that are visually indistinguishable into one class. The dataset contains 100 images for each of the 100 different aircraft model variants. We removed the copyright banner from the images by cutting off the bottom 20 pixels in height. To match the input size of the transformer models, we resized the images to a resolution of 224x224 pixels. We also preprocessed the training images with random crop and horizontal flip layers, similarly to [105]. We partitioned the training set to simulate 65 clients; we set 65 as the number of clients so that, in the IID setting, each client can have at least one per-class (unique) example. At each round, 7 clients out of 65 were randomly selected to participate. Table VI summarizes the dataset-specific configurations.
VIII-C Metrics
We tracked accuracy and KL divergence. Accuracy is measured either on test data or on client $u$’s train data. The KL divergence is computed between the output distribution of the model on client $u$’s data and a uniformly distributed output probability. This metric should work as a proxy for the meaningfulness of model predictions (the greater the divergence the more meaningful the predictions).